OpenStudy (anonymous):

(2p ^{-1})^{3} * (5p ^{2})^{-2}

OpenStudy (anonymous):

(2^3p^(-1*3) ( 5^(-2) p^(2*-2) ) 8p^(-3) x 1/25 * p^(-4) 8/25 * 1/ p^(3+4) = 8/(25p^7)

jimthompson5910 (jim_thompson5910):

\[\large (2p ^{-1})^{3} * (5p ^{2})^{-2}\] \[\large \frac{8}{p^3} * \frac{1}{25p^4}\] \[\large \frac{8}{25p^7}\] So \[\large (2p ^{-1})^{3} * (5p ^{2})^{-2}=\frac{8}{25p^7}\]